This is all one question, I had to split it up in different pictures. please help with all parts, thank you so much in advance!!!

[The following information applies to the questions displayed below.] Shadee Corp. expects to sell 590 sun visors in May andEach visor requires a total of $5.00 in direct materials that includes an adjustable closure that the company purchases fromSuppose that each visor takes 0.90 direct labor hours to produce and Shadee pays its workers $12 per hour. Required: DeterminEach visor requires a total of $5.00 in direct materials that includes an adjustable closure that the company purchases fromEach visor requires a total of $5.00 in direct materials that includes an adjustable closure that the company purchases fromEach visor requires a total of $5.00 in direct materials that includes an adjustable closure that the company purchases from
Answer #1

Solution:

1)

Budgeted Total Sales

May

June

a

Expected Sales Units

590

350

b

Unit Selling Price

$28

$28

a*b

Budgeted Total Sales

$16,520

$9,800

2)

Budgeted Production Units

May

June

a

Expected Sales Unit

590

350

b

Plus: Ending Inventory of visor

40

70

c = a+b

Total needs

630

420

d

Less: Beginning Inventory of visor

75

40

e = c-d

Budgeted Production Units

555

380

3)

Budgeted Cost of Closures Purchases

a

Budgeted Production Units

555.00

380.00

b

Require Raw material closures per unit ($5/1.5 = 3 1/3)

3.33

3.33

a+b=c

Total Closures required

1850.00

1266.67

d

Plus: Ending Closures Unit

18.00

27.00

c-d=e

Total Needs

1868.00

1293.67

f

Less: Beginning Closure Unit

27.00

18.00

e-f=g

Required Purchase of Closure in unit

1841.00

1275.67

h

Cost of Closure per unit

1.50

1.50

g*h

Budgeted Cost of Closures Purchases

2761.50

1913.50

4)

Budeted Manufacturing Overhead

May

June

a

Budgeted Production Units (refer part 2)

555

380

b

Variable Manufacturing Overhead Rate per production unit

$2.25

$2.25

c = a*b

Variable Manufacturing Overhead

$1,248.75

$855.00

d

Fixed Manufacturing Overhead per month

$900

$900

c+d

Total Budgeted Manufacturing Overhead

$2,148.75

$1,755.00

Hope the above calculations, working and explanations are clear to you and help you in understanding the concept of question.... please rate my answer...in case any doubt, post a comment and I will try to resolve the doubt ASAP…thank you

Pls ask separate question for other problems